OpenStudy (anonymous):

solve each equation. state the number and type of roots. PROBLEM: -9x-15=0 how can i get the answer?

OpenStudy (anonymous):

-9x-15=0 -9x=15 x=-15/9

OpenStudy (anonymous):

mhmmm how did you think of that method?

OpenStudy (anonymous):

its just algebra add 15 to both sides, then divide both sides by -9

OpenStudy (anonymous):

okay, i understand that part , but how do i know the type of roots?

OpenStudy (anonymous):

not sure what your asking about with that

OpenStudy (anonymous):

oki here let me give you an example: of what i mean..give me a min

OpenStudy (anonymous):

EXAMPLE: x^(2)+6x+9=0 (x+3)^(2)+0 x+3=0 x=-3 because (x+3) is twice a factor of x^(2)+6x+9, -3 is a double root. thus, the equation has one real repeated root, -3. (FRON THE ALG 2 TEXT BOOK) and i still dont get it ^_^"

OpenStudy (anonymous):

well this equation has no roots because it contains only x^1 so no roots to the equation

OpenStudy (anonymous):

ahhhh oki..... mhm alrighty thanks man ^_~